# Memory Gym: Towards Endless Tasks to Benchmark Memory Capabilities of Agents

### v5

@@ Abstract
Memory Gym presents a suite of 2D partially observable environments, namely Mortar Mayhem, Mystery Path, and Searing Spotlights, designed to benchmark memory capabilities in decision-making agents. These environments, originally with finite tasks, are expanded into innovative, endless formats, mirroring the escalating challenges of cumulative memory games such as “I packed my bag”. This progression in task design shifts the focus from merely assessing sample efficiency to also probing the levels of memory effectiveness in dynamic, prolonged scenarios. To address the gap in available memory-based Deep Reinforcement Learning baselines, we introduce an implementation that integrates Transformer-XL (TrXL) with Proximal Policy Optimization. This approach utilizes TrXL as a form of episodic memory, employing a sliding window technique. Our comparative study between the Gated Recurrent Unit (GRU) and TrXL reveals varied performances across different settings. TrXL, on the finite environments, demonstrates superior sample efficiency in Mystery Path and outperforms in Mortar Mayhem. However, GRU is more efficient on Searing Spotlights. Most notably, in all endless tasks, GRU makes a remarkable resurgence, consistently outperforming TrXL by significant margins.
Website and Source Code: https://github.com/MarcoMeter/endless-memory-gym/
@@ 1.1 Endless Tasks Benchmark Memory Effectiveness and not just Efficiency
To discover the most effective memory approach for complex tasks as mentioned before, insightful benchmarks are needed. However, conventional DRL memory benchmarks, which typically revolve around finite tasks that terminate upon reaching a terminal state regardless of success or failure, fall short in providing the necessary depth. Finite tasks impose a predetermined limit on the demands placed on an agent’s memory. When various memory strategies successfully complete a fixed task, their performance is typically assessed based on efficiency rather than effectiveness. Here, “efficiency” refers to the number of agent-environment interactions required to achieve effective behaviors, commonly known as sample efficiency. Although these methods may reach the same end goal, their true effectiveness and memory capabilities often remain uncertain. For example, a memory mechanism that retains more information or sustains it over extended periods could prove more effective, potentially leading to higher scores, increased returns, better survival times, or improved win rates. This distinction could identify it as a more robust and effective approach compared to others assessed.
Cumulative memory games, such as “I packed my bag” and “Simon”(Morrison and Baer, 1977), overcome the limits of finite tasks. They introduce an automatic curriculum through their inherently progressive and unbounded nature. The game’s expansion depends on the participants’ continuous success. Exploiting the concept of “endlessness” in tasks provide deeper insights into the effectiveness of decision-makers when facing escalating demands of memory retention and recall. These challenges rigorously test an agent’s memory, evaluating both the quantity of information it can retain and the duration for which it can robustly maintain it.
@@ 1.2 Contributions: Novel Memory Benchmark and Transformer-XL Baseline
To exploit the endless behavior of cumulative memory games to thoroughly benchmark memory effectiveness, we enhance our prior work Memory Gym (Pleines et al., 2023), an open-source benchmark, designed to challenge memory-based DRL agents to memorize events across long sequences, generalize, be robust to noise, and be sample efficient. Memory Gym encompasses three distinct finite environments: Mortar Mayhem, Mystery Path, and Searing Spotlights. Each environment presents visual observations, multi-discrete action spaces, and crucially, they cannot be mastered without memory, demanding frequent memory interactions. Building on this foundation, we advance Memory Gym by evolving each environment into an endless format, inspired by the “I packed my bag” game, to further test and develop memory capabilities in DRL agents. standing, no existing DRL memory benchmark possesses such endless tasks.
We further contribute an open-source and easy-to-follow baseline implementation of Transformer-XL (TrXL) (Dai et al., 2019) based on the widely-used DRL algorithm Proximal Policy Optimization (PPO) (Schulman et al., 2017). A recurring theme in the realm of memory-based DRL is the lack of accessible and reproducible baselines. Many significant works, particularly those utilizing transformers (Fortunato et al., 2019; Parisotto et al., 2020; Hill et al., 2021; Lampinen et al., 2021; Parisotto and Salakhutdinov, 2021), have not made their implementations publicly available. As implementation details of DRL algorithms are vital (Engstrom et al., 2020; Andrychowicz et al., 2021; Huang et al., 2022), this lack of transparency poses significant challenges for the research community, hindering progress and reproducibility. Therefore, our accessible implementation can serve as a beacon for the community fostering further research.
In our experimental analysis, we benchmark a TrXL and a recurrent agent, based on the Gated Recurrent Unit (GRU) (Cho et al., 2014), across the original (finite) and the new endless environments of Memory Gym. Our findings indicate that TrXL performs best in terms of sample efficiency on Mystery Path and effectiveness on Mortar Mayhem. However, in the Searing Spotlights environment, GRU demonstrates better sample efficiency. To our greatest surprise, in the endless environments, GRU consistently surpasses TrXL in effectiveness by large margins.
@@ 4 Transformer-XL Baseline
In our experiments, we employ the widely-recognized Deep Reinforcement Learning (DRL) algorithm Proximal Policy Optimization (PPO) (Schulman et al., 2017). To add memory abilities to PPO, either recurrent neural networks (RNN) or attention mechanisms (i.e. transformer) are potential solutions. Prior works have extensively demonstrated the effectiveness of leveraging RNNs, such as Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997) and Gated Recurrent Unit (GRU) (Cho et al., 2014), within the realm of DRL (Mnih et al., 2016; Espeholt et al., 2018; Huang et al., 2022). This also applies to previous studies employing transformers, as Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) and Hierarchical Chunk Attention Memory (HCAM) (Lampinen et al., 2021). However, integrating these memory-enhancing techniques into DRL algorithms is not straightforward, particularly due to unavailable transformer-based DRL implementations. To bridge this gap, we contribute an easy-to-follow baseline implementation of Transformer-XL (TrXL) (Dai et al., 2019) in PPO. Before delving into the transformer-based PPO baseline, we describe the actor-critic model architecture and the losses used during training.
@@ 4.3 Transformer-XL as Memory Mechanism
To begin with, it needs to be clarified what kind of sequential data is fed to a transformer, which is typically composed of multiple stacked layers. Naively, the input data can be conceptualized as a sequence of observations. In the context of visual observations, this is commonly referred to as frame stacking, where consecutive frames are combined. However, instead of expensively stacking raw frames, we leverage the past (i.e. cached) outputs of the model’s embedding module (Figure 4), which serves as the input sequence to the first transformer layer. The subsequent layers make use of the past outputs from their preceding layers. These sequences, those elements we refer to as cached hidden states, are stored in the agent’s episodic memory (Figure 5(a)) and are collected during inference based on the episode’s current timestep $t$. To accommodate computational challenges of potentially endless episodes, we employ a sliding memory window approach, where the input sequence of these hidden states is derived from a fixed-length window. By leveraging cached hidden states in the input sequence, the agent’s memory facilitates the ability to attend to events that extend beyond the boundaries of the fixed-length memory window, aligning with the segment-level recurrence of TrXL (Dai et al., 2019). Note that gradients are not back propagated through the episodic memory because of caching. Once propagated through all transformer layers, a final hidden state is produced to inform the agent’s policy.
@@ 5 Experimental Analysis
We run empirical experiments on Memory Gym using the just introduced TrXL baseline and one based on GRU, both powered by Proximal Policy Optimization (Schulman et al., 2017). The next subsection explains why GRU and TrXL are the only baselines considered. Next, we detail the evaluation protocol, drawing inspiration from Agarwal et al. (2021). Section 5.3 presents results on the finite tasks, showing that TrXL exhibits stronger policies than GRU in Mortar Mayhem, while being more sample efficient in Mystery Path. Searing Spotlights is an exception, where TrXL is less sample efficient. Contradicting our initial expectations and the results on the finite environments, we unveil surprising findings that highlight GRU’s superiority over TrXL in all endless environments. Afterward, we compare the results from the finite and endless environments, demonstrating that the finite settings offer only a limited view of the agents’ true memory capabilities. We then broadly explore several hypotheses on why TrXL might be less effective. Finally, we falsify that recurrent agents are vulnerable to spotlight perturbations when using enhanced hyperparameters (e.g., not normalizing advantage estimates).
@@ 5.2 Evaluation Protocol
A major challenge in DRL research is the statistical uncertainty resulting from high computational costs and the stochastic nature of training algorithms. Due to these expenses, many studies rely on limited data, such as those evaluating PPO, MRA, and HCAM with only three runs per experiment. Under these conditions, Agarwal et al. (2021) propose more robust statistical tools to reliably evaluate the performance of agents trained with DRL. By following their guidelines, we aggregate 5 training runs using the interquartile mean (IQM), which is preferred because the mean is sensitive to outliers and the median requires much more data. To measure and visualize variability, the following sample efficiency curves depict the stratified bootstrapped confidence interval of 0.95 (represented by shaded areas on the plots). To evaluate the generalization ability of our baselines, we extensively test them on 50 novel environment seeds that were excluded from the training process. This evaluation is repeated 3 times for each data point, resulting in a comprehensive assessment based on a total of 750 episodes for every data point presented.
5.3 Finite Environments: Transformer-XL is often more efficient than GRU
(a) Mortar Mayhem Act Grid
(b) Mortar Mayhem Grid
(c) Mortar Mayhem
(d) Mystery Path Grid (Cues On)
(e) Mystery Path Grid
(f) Mystery Path
(g) Searing Spotlights
Figure 6: Performance comparison across instances of Memory Gym’s finite environments. Notably, GRU and TrXL are primarily different in terms of sample efficiency. Only in Mortar Mayhem (c), levels of effectiveness are apparent. Mortar Mayhem Act Grid (a) provides commands as a fully observable vector, bypassing the Clue Task. Both (a) and Mortar Mayhem Grid (b) utilize grid-like locomotion. (c) represents the default Mortar Mayhem task. Mystery Path Grid operates on grid-like locomotion, with (d) rendering the goal and origin to the agent’s observation. (e) and (f) hide these, while (f) is the default Mystery Path task. Searing Spotlights (g) is not varied, while the baselines may leverage observation reconstruction (Obs. Rec.) in this environment.
The introduction of the TrXL baseline motivates conducting benchmarks on the finite environments of Memory Gym, as depicted in Figure 6. Due to the upper bound of the finite tasks, the differences between GRU and TrXL are primarily evident in terms of sample efficiency, with Mortar Mayhem 66(c) being the only exception. We will address this specific scenario after explroing efficiency trends.
Results on Mortar Mayhem (Figures 66(a)-6(c)) show the number of commands that were properly executed during one episode. In Mortar Mayhem Act Grid, GRU and TrXL complete all 10 commands, while TrXL is slightly less efficient. Once the clue task is present (Figure 6(b)), TrXL is more efficient than GRU.
In the Mystery Path environments, we opted not to train agents using observation reconstruction, as the always visible agent is the only entity that requires visual encoding. GRU is only more sample efficient in Mystery Path Grid where the origin and goal are perceivable by the agent (Figure 6(d)). In the other two scenarios (Figures 66(e)-6(f)), TrXL needs less samples to achieve the maximum success rate. If wall-time efficiency is compared, TrXL and GRU are roughly on par (Annex E).
@@ 5.3 Finite Environments: Transformer-XL is often more efficient than GRU
In Searing Spotlights (Figure 6(g)), utilizing the observation reconstruction loss is key. It significantly improves the sample efficiency of GRU and TrXL. The reason for this can be found in rare events: the agent, the coin, and the exit are seldomly visible. Hence, the visual encoder takes advantage from the auxiliary learning signal provided by the observation reconstruction. Nevertheless, both GRU variants, with and without observation reconstruction, exhibit notably higher sample efficiency compared to their respective TrXL counterparts.
Returning to the exception noted in Mortar Mayhem (Figure 66(c)), we quantify effectiveness levels only for GRU, GRU with observation reconstruction, and TrXL. GRU completes 6 commands, observation reconstruction reduces GRU’s performance to 5, while TrXL without this auxiliary head performs poorly, completing only 1 command. Conversely, TrXL with observation reconstruction effectively resolves all 10 commands, though its true limit remains uncertain. As previously argued in Section 3.1, increasing the difficulty is necessary to accurately assess TrXL’s limits, which we will further investigate by exploiting the automatic curriculum of the endless environments.
@@ 5.4 Endless Environments: GRU is more effective than Transformer-XL
(a) Endless Mortar Mayhem
(b) Endless Mystery Path
(c) Endless Searing Spotlights
Figure 7: GRU’s consistent superiority over TrXL is revealed in Memory Gym’s endless environments. The less opaque line marked with crosses depicts the interquartile mean (IQM) episode length that refers to the right Y-Axis.
@@
Notably, the most substantial gap between GRU and TrXL emerges in the results of Endless Mortar Mayhem (Figure 7(a)). Within this context, GRU attains an impressive IQM of 125 executed commands using roughly 300 million steps, whereas TrXL only attains an IQM of 18 commands. The incorporation of observation reconstruction exacerbates this performance difference, leading TrXL to a further deterioration, resulting in an IQM of 10. GRU exhibits consistently longer IQM episode lengths, with TrXL achieving an IQM of 288 steps and GRU reaching a peak IQM episode length of 3462 steps.
The outcomes depicted in Figure 7(b) illustrate the results obtained from the Endless Mystery Path environment. In this case, GRU remains more effective than TrXL, although the gap between their performances is narrower compared to Endless Mortar Mayhem. GRU achieves an IQM undiscounted return of 2.65, with TrXL achieving 2.02. Although neither memory approach converges within the allocated training budget of 819 million steps, GRU displays a more pronounced upward trend in performance. In terms of episode length, neither agent achieves lengths as seen in Endless Mortar Mayhem. GRU attains an IQM episode length of 600 steps in Endless Mystery Path, while TrXL reaches up to 400 steps. The divergence between the returns and the episode lengths are attributed to the episode’s dynamics, where agents tend to fall off after significant progress, necessitating more steps to regain lost ground and accumulate further rewards.
In Endless Searing Spotlights (Figure 7(c)), GRU is about 3.3 times more effective than TrXL. Both agents utilize the observation reconstruction loss, but TrXL converges to an IQM of 6 collected coins after approximately 200 million steps, while GRU achieves an IQM of about 20 coins, continuing to show an upward trend in its performance. The IQM episode length for TrXL is 200 steps, while GRU achieves a longer IQM episode length of 600 steps.
@@
In this subsection, we highlight that the insights derived from finite environments provide only a truncated view of the true memory capabilities of the agents benchmarked. To recall, higher memory effectiveness implies the ability to memorize more information over extended periods.
Mortar Mayhem stands as an exception among finite environments, where varied levels of effectiveness are observable. One might assume that these differences would align with the results from endless tasks, yet this alignment does not occur. Notably, in the endless setting, GRU completes significantly more commands than those 6 observed in the finite environment. This increased quantity can be attributed to two key factors. Firstly, the environment’s curriculum begins at an easier level by only demanding one command initially. Secondly, Endless Mortar Mayhem allows for two strategies for memorizing commands: the direct visualization of the command and the agent’s ability to infer commands based on its past actions. This may allow for the agent’s policy to adapt faster to growing difficulties.
In the Mystery Path environment, challenges are limited to a maximum path length of 17 tiles, providing only a basic indication of the memory requirements. This measure is approximate, as the procedurally generated path may include multiple turns, increasing the likelihood of the agent falling off. Both GRU and TrXL effectively master this challenge, but they actually surpass these capabilities in the endless version. The undiscounted return reflects the length of the path to the farthest point reached by the agent: 26.5 tiles for GRU and 20 tiles for TrXL, indicating that both models are capable of retaining and recalling more information than demonstrated in the finite setting. As scores correlate with episode duration, agents must maintain robust memory over prolonged periods.
Quantifying performance in Searing Spotlights involves navigating to two entities (coin and exit), while the agent must maintain its hidden position. In finite tasks, effective episodes typically last fewer than 50 steps. However, in Endless Searing Spotlights, GRU and TrXL can collect 20 and 6 coins and maintain their hidden positions for 600 and 200 steps, respectively. This performance demonstrates that both agents not only achieve higher scores but also sustain their concealed positions for longer durations than observed in the finite version of Searing Spotlights.
To conclude, endless tasks reveal more effective memory capabilities in both baselines beyond what is evident in finite tasks.

### v6

@@ Abstract
Memory Gym presents a suite of 2D partially observable environments, namely Mortar Mayhem, Mystery Path, and Searing Spotlights, designed to benchmark memory capabilities in decision-making agents. These environments, originally with finite tasks, are expanded into innovative, endless formats, mirroring the escalating challenges of cumulative memory games such as “I packed my bag”. This progression in task design shifts the focus from merely assessing sample efficiency to also probing the levels of memory effectiveness in dynamic, prolonged scenarios. To address the gap in available memory-based Deep Reinforcement Learning baselines, we introduce an implementation within the open-source CleanRL library that integrates Transformer-XL (TrXL) with Proximal Policy Optimization. This approach utilizes TrXL as a form of episodic memory, employing a sliding window technique. Our comparative study between the Gated Recurrent Unit (GRU) and TrXL reveals varied performances across our finite and endless tasks. TrXL, on the finite environments, demonstrates superior effectiveness over GRU, but only when utilizing an auxiliary loss to reconstruct observations. Notably, GRU makes a remarkable resurgence in all endless tasks, consistently outperforming TrXL by significant margins.
Website and Source Code: https://marcometer.github.io/jmlr_2024.github.io/
@@ 1.1 Endless Tasks Benchmark Memory Effectiveness and not just Efficiency
To discover the most effective memory approach for complex tasks, as mentioned earlier, insightful benchmarks are needed. However, conventional DRL memory benchmarks, which typically revolve around finite tasks that terminate upon reaching a terminal state regardless of success or failure, fall short in providing the necessary depth. Finite tasks impose a predetermined limit on the demands placed on an agent’s memory. When various memory strategies successfully complete a fixed task, their performance is typically assessed based on efficiency rather than effectiveness. Here, “efficiency” refers to the number of agent-environment interactions required to achieve effective behaviors, commonly known as sample efficiency. Although these methods may reach the same end goal, their true effectiveness and memory capabilities often remain uncertain. For example, a memory mechanism that retains more information or sustains it over extended periods could prove more effective, potentially leading to higher scores, increased returns, better survival times, or improved win rates. This distinction could identify it as a more robust and effective approach compared to others assessed.
Cumulative memory games, such as “I Packed My Bag” and “Simon”(Morrison and Baer, 1977), overcome the limitations of finite tasks by introducing an automatic curriculum through their inherently progressive and unbounded nature. These games represent a special instance of endlessness: while theoretically offering a truly endless challenge, in practice, episodes inevitably terminate due to the finite nature of memory. The game’s expansion relies on the agents’ continuous success, thereby embodying the concept of “endlessness” in tasks. This approach provides deeper insights into the effectiveness of decision-makers when facing escalating demands for memory retention and recall. Consequently, these challenges rigorously test an agent’s memory, evaluating both the quantity of information it can retain and the duration for which it can robustly maintain it.
@@ 1.2 Contributions: Novel Memory Benchmark and Transformer-XL Baseline
To exploit the endless behavior of cumulative memory games to thoroughly benchmark memory effectiveness, we enhance our prior work Memory Gym (Pleines et al., 2023), an open-source benchmark, designed to challenge memory-based DRL agents to memorize events across long sequences, generalize, be robust to noise, and be sample efficient. Memory Gym encompasses three distinct finite environments: Mortar Mayhem, Mystery Path, and Searing Spotlights. Each environment presents visual observations, multi-discrete action spaces, and crucially, they cannot be mastered without memory, demanding frequent memory interactions. Building on this foundation, we advance Memory Gym by evolving each environment into an endless format, inspired by the “I packed my bag” game, to further test and develop memory capabilities in DRL agents. To our understanding, no existing DRL memory benchmark possesses such endless tasks.
We further contribute an easy-to-follow baseline implementation of Transformer-XL (TrXL) (Dai et al., 2019) to the open-source CleanRL library (Huang et al., 2022b), based on the widely used Deep Reinforcement Learning (DRL) algorithm Proximal Policy Optimization (PPO) (Schulman et al., 2017). A recurring theme in the realm of memory-based DRL is the lack of accessible and reproducible baselines. Many significant works, particularly those utilizing transformers (Fortunato et al., 2019; Parisotto et al., 2020; Hill et al., 2021; Lampinen et al., 2021; Parisotto and Salakhutdinov, 2021), have not made their implementations publicly available. As implementation details of DRL algorithms are vital (Engstrom et al., 2020; Andrychowicz et al., 2021; Huang et al., 2022a), this lack of transparency poses significant challenges for the research community, hindering progress and reproducibility. Therefore, our accessible implementation can serve as a beacon for the community fostering further research.
In our experimental analysis, we benchmark a TrXL and a recurrent agent, based on the Gated Recurrent Unit (GRU) (Cho et al., 2014), across the original (finite) and the new endless environments of Memory Gym. Our findings indicate that TrXL, only with the help of observation reconstruction as auxiliary loss, is most effective on the finite environments. To our greatest surprise, in the endless environments, GRU consistently surpasses TrXL by large margins.
@@ 4 Transformer-XL Baseline
In our experiments, we employ the widely-recognized Deep Reinforcement Learning (DRL) algorithm Proximal Policy Optimization (PPO) (Schulman et al., 2017). To add memory abilities to PPO, either recurrent neural networks (RNN) or attention mechanisms (i.e. transformer) are potential solutions. Prior works have extensively demonstrated the effectiveness of leveraging RNNs, such as Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997) and Gated Recurrent Unit (GRU) (Cho et al., 2014), within the realm of DRL (Mnih et al., 2016; Espeholt et al., 2018; Huang et al., 2022a). This also applies to previous studies employing transformers, as Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) and Hierarchical Chunk Attention Memory (HCAM) (Lampinen et al., 2021). However, integrating these memory-enhancing techniques into DRL algorithms is not straightforward, particularly due to unavailable transformer-based DRL implementations. To bridge this gap, we contribute an easy-to-follow baseline implementation of Transformer-XL (TrXL) (Dai et al., 2019) in PPO. Before delving into the transformer-based PPO baseline, we describe the actor-critic model architecture and the losses used during training.
@@ 4.3 Transformer-XL as Memory Mechanism
To begin with, it needs to be clarified what kind of sequential data is fed to a transformer, which is typically composed of multiple stacked layers. Naively, the input data can be conceptualized as a sequence of observations. In the context of visual observations, this is commonly referred to as frame stacking, where consecutive frames are combined. However, instead of expensively stacking raw frames, we leverage the past (i.e. cached) outputs of the model’s embedding module (Figure 4), which serves as the input sequence to the first transformer layer. The subsequent layers make use of the past outputs from their preceding layers. These sequences, those elements we refer to as cached hidden states, are stored in the agent’s episodic memory (Figure 5(a)) and are collected during inference based on the episode’s current timestep $t$. To accommodate computational challenges of potentially endless episodes, we employ a sliding memory window approach, where the input sequence of these hidden states is derived from a fixed-length window. By leveraging cached hidden states in the input sequence, the agent’s memory facilitates the ability to attend to events that extend beyond the boundaries of the fixed-length memory window, aligning with the segment-level recurrence of TrXL (Dai et al., 2019). Let $N$ represent the number of layers and $L$ denote the window length. The maximum context length $L_{\text{max}}$ is given by
$L_{\text{max}}=N\times(L-1)+1\text{.}$ (7)
Note that gradients are not back propagated through the episodic memory because of caching. Once propagated through all transformer layers, a final hidden state is produced to inform the agent’s policy.
@@ 5 Experimental Analysis
We run empirical experiments on Memory Gym using the just introduced TrXL baseline and one based on GRU, both powered by Proximal Policy Optimization (Schulman et al., 2017). The next subsection explains why GRU and TrXL are the only baselines considered. Next, we detail the evaluation protocol. Section 5.3 presents results on the finite tasks, showing that TrXL exhibits stronger policies than GRU, but only when leveraging observation reconstruction as an auxiliary loss. Contradicting our initial expectations and the results on the finite environments, we unveil surprising findings that highlight GRU’s superiority over TrXL in all endless environments. Afterward, we compare the results from the finite and endless environments, demonstrating that the finite settings offer only a limited view of the agents’ true memory capabilities. We then broadly explore several hypotheses on why TrXL might be less effective. Finally, we falsify that recurrent agents are vulnerable to spotlight perturbations when using enhanced hyperparameters (e.g., not normalizing advantage estimates).
@@ 5.2 Evaluation Protocol
A major challenge in DRL research is the statistical uncertainty resulting from high computational costs and the stochastic nature of training algorithms (Agarwal et al., 2021; Jordan et al., 2024). Due to these expenses, many studies rely on limited samples, such as those evaluating PPO, MRA, and HCAM with only three runs (i.e., seeds) per experiment. While our empirical protocol does not fully resolve this issue, it mitigates it by repeating each experiment five times on independent seeds. The code and data behind all plots are publicly available at https://marcometer.github.io/jmlr_2024.github.io/.
The forthcoming results on the finite and endless environments (Sections 5.3 and 5.4) display the mean performance of each individual run. Instead of aggregating these runs, we take advantage of the available space to provide denser information. Each data point of a single run is an average of 50 episode seeds, repeated three times to account for the agent’s stochastic policy. Although five samples may still be insufficient for drawing strong conclusions, we increase reliability by aggregating the trained agents’ performances across both finite and endless environments. These aggregation plots (Figures 6(d) and 7(d)) illustrate the mean and standard deviation of 15 independent seeds, allowing for a more robust statistical analysis. The remaining sample efficiency curves, which are used for supplementary experiments (e.g., exploring TrXL’s low effectiveness on endless tasks), average five runs and display standard deviations.
5.3 Finite Environments: Transformer-XL Relies on Obs. Reconstruction
(a) Mortar Mayhem
(b) Mystery Path
(c) Searing Spotlights
(d) Aggregation and Normalization Across Finite Tasks
Figure 6: Performances across Memory Gym’s finite environments. For each agent, the mean performance of 5 independent seeds is shown (a, b, and c). (d) shows the mean performance and the standard deviation when normalizing and aggregating across the finite tasks.
The introduction of the TrXL baseline motivates conducting benchmarks on the finite environments of Memory Gym, as depicted in Figure 6. Instead of comparing the success rates on the full tasks, we consider a more fine-grained task progression that provides denser information.
In Mortar Mayhem (Figure 66(a)), the task is completed when an agent successfully executes 10 commands. The only agent that completes the entire task is the TrXL agent, which benefits from observation reconstruction. The GRU agent, which does not use this auxiliary loss, achieves an average completion of 6.2 commands at best and 4.2 commands at worst. TrXL without reconstruction ranges from 1.3 to 5 commands.
In Mystery Path, task progress is based on the proportion of the path completed. The path lengths of the evaluated episodes vary between 7 and 14 tiles. In this environment, all agents are effective, with the TrXL agents being more sample efficient than their GRU counterparts. However, when wall-time efficiency is compared, TrXL and GRU are about on par because TrXL is slower (Annex E).
@@
In Searing Spotlights (Figure 6(c)), the reported task progression measures 50% for collecting the coin, with the remainder dedicated to successfully terminating the episode by using the exit. The results demonstrate that utilizing the observation reconstruction loss is key, significantly improving the effectiveness of TrXL and the sample efficiency of GRU. This improvement is due to rare events where the agent, the coin, and the exit are seldomly visible. Thus, the visual encoder benefits from the auxiliary learning signal provided by the observation reconstruction. Nevertheless, both GRU variants, exhibit notably higher sample efficiency compared to TrXL.
As a final assessment, we aggregate and average the normalized task progression in Figure 66(d). Each curve represents the mean and standard deviation of 15 independent runs. This statistical measure demonstrates that TrXL relies on observation reconstruction to complete 99% of the tasks on average. Without it, its mean task progression drops to 56%. GRU without reconstruction loss reaches a task progression of 83%. When leveraging observation reconstruction, GRU appears more sample efficient, with a mean task completion of 82%.
@@ 5.4 Endless Environments: GRU is more effective than Transformer-XL
(a) Mortar Mayhem
(b) Mystery Path
(c) Searing Spotlights
(d) Aggregation and Normalization Across Finite Tasks
Figure 7: GRU’s consistent superiority over TrXL is revealed in Memory Gym’s endless environments. For each agent, the mean performance of 5 independent seeds is shown (a, b, and c). (d) shows the mean performance and the standard deviation when normalizing and aggregating across the endless tasks.
@@
Notably, the most substantial gap between GRU and TrXL emerges in the results of Endless Mortar Mayhem (Figure 7(a)). In this context, GRU attains an impressive range of 84 to 120 executed commands using roughly 300 million steps, whereas both TrXL baselines only attain an average of 17 commands. The best data point for the recurrent agent measures 140 successfully executed commands on average for a single training run. When taking the mean of the 5 GRU runs, the best point in time scores 115 commands with a standard deviation of 15.
The outcomes depicted in Figure 7(b) illustrate the results obtained from the Endless Mystery Path environment. In this case, GRU remains more effective than TrXL, although the gap between their performances is narrower compared to Endless Mortar Mayhem. In this setting, the best TrXL run visits 21 tiles of the path, while GRU visits 29 tiles at best. After roughly 150 million steps, all GRU curves continuously distance themselves from TrXL. Upon training completion, the mean of all GRU runs is 26 tiles, while the TrXL ones average 20 tiles.
In Endless Searing Spotlights (Figure 7(c)), both agents utilize the observation reconstruction loss. TrXL is capable of collecting 7.5 coins at maximum, while GRU’s best performance is nearly 4 times higher with 29.8 collected coins. The final performance after training measures 23.9 coins for GRU and 6.7 ones for TrXL.
All results obtained from the endless environments show a clear performance gap between GRU and TrXL. This is also salient in the aggregation plot of Figure 7(d), which combines all GRU and TrXL runs, including those from Endless Searing Spotlights that use observation reconstruction. Hence, both the green and red curves aggregate 15 independent seeds. The data is normalized by the maximum mean score of the GRU agent in each environment. Figure 7(d) shows that the GRU agent peaks at 0.98, while TrXL reaches a normalized score of 0.39. Additionally, the error bars of the standard deviations do not overlap after roughly 250 million steps.
@@
In this subsection, we highlight that the insights derived from finite environments provide only a truncated view of the true memory capabilities of the benchmarked agents. To recall, higher memory effectiveness implies the ability to memorize more information over extended periods.
Mortar Mayhem stands as an exception among finite environments, where varied levels of effectiveness are observable. One might assume that these differences would align with the results from endless tasks, yet this alignment does not occur. Notably, in the endless setting, GRU completes significantly more commands than the 6.2 observed in the finite environment. This increased quantity may be attributed to two key factors. Firstly, the environment’s curriculum begins at an easier level by only demanding one command initially. Secondly, Endless Mortar Mayhem allows for two strategies for memorizing commands: the direct visualization of the command and the agent’s ability to infer commands based on its past actions. This may allow for the agent’s policy to adapt faster to growing difficulties. Regarding episode lengths, when completing 10 commands in the finite tasks, episodes last for 274 steps. In the endless task, GRU plays for more than 3000 steps and TrXL for more than 400 steps.
In the Mystery Path environment, challenges are limited to a maximum path length of 14 tiles, providing only a basic indication of the memory requirements. This measure is approximate, as the procedurally generated path may include multiple turns, increasing the likelihood of the agent falling off the path. Both GRU and TrXL effectively master this challenge, but they surpass these capabilities in the endless version. While successful episodes in the finite task last for about 100 steps, GRU reaches more than 600 steps on average in the endless task, while TrXL approaches 400 steps.
Quantifying performance in Searing Spotlights involves navigating to two entities (coin and exit), while the agent must maintain its hidden position. In this finite task, effective episodes typically last fewer than 50 steps. However, in Endless Searing Spotlights, GRU and TrXL can gather more coins and maintain their hidden positions for 691 and 240 steps on average, respectively. This performance demonstrates that both agents not only achieve higher scores but also sustain their concealed positions for longer durations than observed in the finite version of Searing Spotlights.
To conclude, endless tasks demonstrate their potential to reveal more effective memory capabilities in both baselines beyond what is evident in finite tasks, as more information needs to be memorized and maintained over prolonged time horizons.
To conclude, finite environments provide only a truncated view of the agents’ capabilities. Instead, the endless environments demonstrate their potential to reveal more effective memory capabilities in both baselines beyond what is evident in finite tasks, as more information needs to be memorized and maintained over ever-growing time horizons.